# MegaScale-MoE: Large-Scale Communication-Efficient Training of Mixture-of-Experts Models in Production

### v2

@@ Abstract
Xin Jin, Xin Liu
1 Introduction
As the size of Large Language Models (LLMs) [7, 47, 16] grow, so does the scale of their training regimes. The escalation in training scale has made efficiency improvements not just desirable but crucial [17]. As a company building AI products for billions of users, we remain committed to training LLMs with hundreds of billions of parameters on thousands of GPUs. Consequently, even marginal gains in training efficiency can significantly reduce computational resource consumption and training time, directly influencing the feasibility and sustainability of developing state-of-the-art LLMs.
Within the landscape of LLM architectures, Mixture-of-Experts (MoE) models stand out for their sparse activation [7, 16, 9, 44], which dynamically routes input tokens to a selected set of specialized network components, known as experts, rather than to all parameters. This design leads to sub-linear scaling of FLOPs required as the model size increases, thereby significantly reducing the computational cost. Recent industrial advancements [8, 40, 2, 3, 25] have demonstrated the potential of MoE models, achieving an order-of-magnitude reduction in training cost compared to dense models with equivalent model quality.
@@ 1 Introduction
Moreover, as hardware advances, the imbalance between computation and communication becomes increasingly pronounced, with communication overhead growing more dominant. Alongside improvements in model architectures, hardware capabilities have evolved rapidly, with GPUs achieving significantly higher processing speeds (Figure 1). Concurrently, reductions in training precision have been adopted to enhance efficient and cost-effective training [36, 25]. These trends lead to a scenario where the raw computation time decreases, making the relative impact of communication overhead a more critical bottleneck. For instance, simply extending existing tensor parallelism to multi-node setups has been observed to push communication overhead beyond 50% in certain cases. As a result, optimizing communication is essential for sustaining and improving the scalability of MoE model training, particularly in distributed environments where frequent data synchronization across multiple GPUs is required.
In this paper, we present the design, implementation, and operational experience of MegaScale-MoE, a production system optimized for efficient large-scale MoE training. By meticulously addressing the communication bottleneck, MegaScale-MoE strives to push the boundaries of MoE training, achieving significant improvements in performance and efficiency.
MegaScale-MoE addresses the communication problem in MoE training from three key aspects. First, MegaScale-MoE reduces the communication volume by customizing parallelism strategies for the attention and FFN modules in each MoE layer. We compare the parallelism strategies in existing LLM training frameworks, comprehensively considering their impact on large-scale training, including the communication volume and whether communication can be effectively overlapped (i.e., whether it lies on the critical path). Based on this analysis, we select the optimal combination of parallelism strategies for MoE training.
@@
MegaScale-MoE is deployed in our datacenters to train MoE models for our products. Compared to the state-of-the-art open-source LLM training framework, Megatron-LM [46], MegaScale-MoE achieves up to 1.88 higher MFU (Model FLOPs Utilization) when training a 352B MoE model on 1,440 NVIDIA Hopper GPUs. With comprehensive communication optimizations, MegaScale-MoE powers large-scale training in our production, efficiently scaling to trillions of parameters and thousands of GPUs while saving millions of GPU hours.
2 Background
2.1 Mixture-of-Experts for Transformer
@@
The Mixture of Experts (MoE) mechanism is an advanced approach designed to boost the performance of Transformer [49] models, which are increasingly pivotal in the realm of LLMs [16, 7, 2, 25]. It extends the Transformer architecture by integrating multiple expert networks within the feed-forward network (FFN) component. As illustrated in Figure 2, MoE models dynamically route input tokens to the most relevant experts based on their characteristics. This routing is managed by a trainable gating mechanism that selects the best-suited experts for each token. This architectural innovation enables MoE models to scale in capacity without a proportional increase in inference costs, as only a subset of experts is activated for each input.
2.2 Large-scale LLM Training
@@
Training large language models at scale on tens of thousands of GPUs is a complex system engineering challenge that requires multiple systems techniques. To distribute the training workload, a combination of parallelism strategies such as data, tensor, and pipeline parallelism is necessary [46, 41, 17], as each approach has limitations that prevent relying on a single method for effective scaling.
@@
Data parallelism uniformly distributes the training data across all devices, with each device replicating the model parameters and optimizer states. To synchronize the parameters after each training iteration, data parallelism performs an all-reduce communication operation. Zero Redundancy Optimizer (ZeRO) [38] improves over data parallelism by distributing model states across all participating devices. ZeRO unfolds across three progressive stages, each designed to increasingly conserve memory, though this comes with the trade-off of elevated communication.
@@
Tensor parallelism distributes compute-intensive tensor operations over multiple devices, enabling parallel computation and significantly accelerating the training process. The specific partitioning strategy and the dependencies among operators within the model dictate that tensor parallelism may necessitate gathering split inputs (all-gather) or merging outputs (reduce-scatter). In LLM training, operators like LayerNorm and Dropout, though less compute-intensive, require substantial activation memory. To tackle this problem, a variant of tensor parallelism known as sequence parallelism [18] is proposed, which partitions these operators along the dimension of sequence length. For long-context training, several works [46, 14, 1] apply sequence parallelism or tensor parallelism to different operators in self-attention. Figure 3 illustrates the mainstream parallelism strategies for attention, namely tensor, sequence, and context parallelism (TP, SP, and CP), which we analyze in §3.1.
@@
Pipeline parallelism enhances efficiency by dividing model layers into stages that are processed on different devices, enabling pipelined execution. Each batch is split into several micro-batches for this purpose. To minimize pipeline bubbles, various scheduling strategies have been developed, e.g., GPipe [12], PipeDream 1F1B [31] and Interleaved 1F1B [32], etc. Megatron-LM adopts Interleaved 1F1B pipeline scheduling, further dividing each stage on one device into multiple virtual stages to reduce the pipeline bubble rate.
@@
3 Communication-Efficient Parallelism
@@
Figure 4 shows the design space of parallelism strategies for large-scale MoE training, excluding the outermost data parallelism. We start with inter-node parallelism. Expert parallelism alleviates memory pressure from MoE models’ large parameter size by distributing experts across nodes but incurs per-layer cross-node communication, harming training efficiency. Similarly, tensor parallelism’s high communication overhead makes it more efficient to limit TP to a single node. Following prior work [17], we adopt pipeline parallelism to distribute model parameters, reduce communication, and overlap communication of different micro-batches.
Prior large-scale MoE training systems, such as Megatron-LM [46] and DeepSpeed-MoE [40], incorporate tensor parallelism to scale up training by partitioning the model parameters within the node. However, in our practice, we observe two issues with this approach: (1) TP partitions the expert dimension, which negatively impacts GEMM efficiency; and (2) TP introduces significant communication overhead, which remains constant as the parallelism size increases, eventually causing communication to exceed computation on modern hardware.
@@
3.1 Sequence Parallelism for Attention
Due to the inherent parallelizability of the expert components in MoE models, most prior work on MoE training [40, 20] focuses on optimizing expert parallelism, while data parallelism (DP) is typically applied to the non-MoE components such as attention. However, when scaling up MoE training, this approach proves insufficient due to the activation memory consumption. This issue arises because DP splits the batch dimension both across and within nodes. Compared to other intra-node parallelism strategies shown in Figure 4, applying DP to attention forces each GPU within a node to process one micro-batch simultaneously, increasing the activation size by 8, which often results in out-of-memory issues.
@@ 3.1 Sequence Parallelism for Attention
We adopt sequence parallelism (SP), as proposed in DeepSpeed-Ulysses [14], to scale MoE training and effectively reduce communication along the critical path. SP is commonly used in long-context training to address memory challenges associated with long inputs. We find it also works well in large-scale MoE training. First, it significantly reduces communication overhead compared to TP, especially when using grouped-query attention [4]. Second, while it introduces some parameter redundancy and increased communication overhead during parameter synchronization, the unique characteristics of MoE models make these trade-offs manageable and acceptable.
@@
where represents the ratio between the number of query heads and that of key-value heads. Assuming the model is trained on a NVIDIA Hopper GPU workstation with an NVLink domain of size 8, the communication latency for sequence parallel attention can be reduced to about one-fourth of that required by tensor-parallel attention.
Data communication & memory overhead. A notable difference between SP and TP attention is how parameters are distributed across devices: TP shards the attention weights, while SP replicates them. This raises the concern about the potential increase in communication overhead for synchronizing gradients and parameters. Counterintuitively, given the intra- and inter-node bandwidth asymmetry and the adoption of hierarchical communication operations in modern communication libraries [33] as shown in Figure 5 and analyzed in Appendix A.1, although SP attention requires synchronization of more parameters compared to TP attention, the difference in communication overhead is minimal in practical scenarios.
@@
Balanced vs. imbalanced. In addition to the Ulysses-style SP attention, we also explored other forms, including context parallelism (CP) [1], which partitions all activations along the sequence dimension. CP attention, however, faces workload imbalance due to causal masking in attention, as each token only attends to previous tokens. To mitigate this, we attempted the zigzag strategy by grouping the head and tail partitions of the sequence on the same GPU, although achieving perfect balance remains challenging. Consequently, in large-scale training, the entire training process is often constrained by the most imbalanced data batch. Moreover, this imbalance disturbs the training pipeline, thereby reducing overall training efficiency.
3.2 Expert Parallelism for Feed-forward Network
@@
In practical training, all-to-all communication is less efficient than all-gather and reduce-scatter, as it requires each worker to communicate with all others, whereas all-gather and reduce-scatter follow a ring-based communication pattern with only neighboring workers. As shown in Figure 7, the communication time for these three operations in Mixtral-87B reveals that when top- > 6, the all-gather-based EP implementation is more efficient.
@@
Load balance. A well-known challenge in MoE model training is load balancing across experts [20, 24]. To address this, we use auxiliary loss and token dropping to balance the workload across GPUs within each node. Similar to DeepSeek-V2 [24], we treat the experts placed on the same GPU as a group and calculate the balance loss and computational capacity for each device rather than for each individual expert.
4 Communication-computation Overlap
@@
4.1 Inter-operator Overlap
@@
Selective activation rematerialization. The holistic scheduling strategy also helps reduce memory usage without compromising training speed. Compared to dense models with equivalent computational requirements, MoE models exert significantly higher memory pressure during training due to their parameter count being several times larger. In addition to employing ZeRO optimizations [38] to eliminate redundant optimizer states across DP groups, we further optimize memory usage through selective activation rematerialization. This approach reduces activation memory requirements by re-performing computation and communication operators that can be overlapped with other necessary operators.
Figure 8a illustrates the forward pass of a Mixtral [16] MoE layer and highlights key activations produced during this process. MegaScale-MoE strategically retains activations that are computationally expensive to recompute, while recalculating others generated by memory-intensive operations or communication operations. This minimizes dependencies on backward computation, enabling rematerialization operations to overlap with other computations and communications, avoiding delays in the critical path. For example, as shown in Figure 8b, the backward pass of the GroupedGEMM operator for FC2 requires the activation fc2_in and the gradient of fc2_out (denoted as fc2_out) as inputs. MegaScale-MoE recomputes fc2_in and overlaps this operator with gradient communication (i.e., all-gather for ffn_out). Similarly, ffn_in is obtained through re-performing RMSNorm and all-gather, with these operators hidden within the preceding communication and the FC2 GroupedGEMM, respectively. MegaScale-MoE also places the weighted sum of ffn_out immediately after the SwiGLU [43] activation function to eliminate the need to store ffn_out. This reordering ensures computational consistency by avoiding operators that cross non-linear boundaries.
4.2 Intra-operator Overlap

### v3

@@ Abstract.
ACM Reference Format:
Chao Jin, Ziheng Jiang, Zhihao Bai, Zheng Zhong, Juncai Liu, Xiang Li, Ningxin Zheng, Xi Wang, Cong Xie, Qi Huang, Wen Heng, Yiyuan Ma, Wenlei Bao, Size Zheng, Yanghua Peng, Haibin Lin, Xuanzhe Liu, Xin Jin, Xin Liu. 2025. MegaScale-MoE: Large-Scale Communication-Efficient Training of Mixture-of-Experts Models in Production. In EuroSys ’26, April 27–30, 2026, Edinburgh, UK. In ACM, New York, NY, USA, 17 pages. https://doi.org/10.1145/3767295.3769325
1. Introduction
As the size of Large Language Models (LLMs) (Chowdhery et al., 2023; Touvron et al., 2023; Jiang et al., 2024b) grow, so does the scale of their training regimes. The escalation in training scale has made efficiency improvements not just desirable but crucial (Jiang et al., 2024a). As a company building AI products for billions of users, we remain committed to training LLMs with hundreds of billions of parameters on thousands of GPUs. Consequently, even marginal gains in training efficiency can significantly reduce computational resource consumption and training time, directly influencing the feasibility and sustainability of developing state-of-the-art LLMs.
Within the landscape of LLM architectures, Mixture-of-Experts (MoE) models stand out for their sparse activation (Chowdhery et al., 2023; Jiang et al., 2024b; Fedus et al., 2022; Shazeer et al., 2017), which dynamically routes input tokens to a selected set of specialized network components, known as experts, rather than to all parameters. This design leads to sub-linear scaling of FLOPs required as the model size increases, thereby significantly reducing the computational cost. Recent industrial advancements (Du et al., 2022; Rajbhandari et al., 2022; dbr, 2025; gro, 2025; Liu et al., 2024b) have demonstrated the potential of MoE models, achieving an order-of-magnitude reduction in training cost compared to dense models with equivalent model quality.
@@ 1. Introduction
Moreover, as hardware advances, the imbalance between computation and communication becomes increasingly pronounced, with communication overhead growing more dominant. Alongside improvements in model architectures, hardware capabilities have evolved rapidly, with GPUs achieving significantly higher processing speeds (Figure 1). Concurrently, reductions in training precision have been adopted to enhance efficient and cost-effective training (Peng et al., 2023; Liu et al., 2024b). These trends lead to a scenario where the raw computation time decreases, making the relative impact of communication overhead a more critical bottleneck. For instance, simply extending existing tensor parallelism to multi-node setups has been observed to push communication overhead beyond 50% in certain cases. As a result, optimizing communication is essential for sustaining and improving the scalability of MoE model training, particularly in distributed environments where frequent data synchronization across multiple GPUs is required.
In this paper, we present the design, implementation, and operational experience of MegaScale-MoE, a production system optimized for efficient large-scale MoE training. By meticulously addressing the communication bottleneck, MegaScale-MoE strives to push the boundaries of MoE training, achieving significant improvements in performance and efficiency. Based on the insight that the key architectural distinctions between MoE and dense models are intra-layer, which is the primary source of the communication overhead, MegaScale-MoE confines each MoE layer to within a single node, utilizing high-bandwidth NVLink. Our analysis (§3) and evaluation (§6) show that despite the cross-node expert parallelism common in existing systems (Liu et al., 2024b; Hwang et al., 2023), our approach effectively scales MoE training to models of several hundred billion parameters on thousands of GPUs.
Specifically, MegaScale-MoE addresses the communication problem in MoE training from three key aspects. First, MegaScale-MoE reduces the communication volume by customizing parallelism strategies for the attention and FFN modules in each MoE layer. We compare the parallelism strategies in existing LLM training frameworks, comprehensively considering their impact on large-scale training, including the communication volume and whether communication can be effectively overlapped (i.e., whether it lies on the critical path). Based on this analysis, we select the optimal combination of parallelism strategies for MoE training.
@@
MegaScale-MoE is deployed in our datacenters to train MoE models for our products. Compared to the state-of-the-art open-source LLM training framework, Megatron-LM (Shoeybi et al., 2019), MegaScale-MoE achieves up to 1.88 higher MFU (Model FLOPs Utilization) when training a 352B MoE model on 1,440 NVIDIA Hopper GPUs. With comprehensive communication optimizations, MegaScale-MoE powers large-scale training in our production, efficiently scaling to trillions of parameters and thousands of GPUs while saving millions of GPU hours.
2. Background
2.1. Mixture-of-Experts for Transformer
@@
The Mixture of Experts (MoE) mechanism is an advanced approach designed to boost the performance of Transformer (Vaswani et al., 2017) models, which are increasingly pivotal in the realm of LLMs (Jiang et al., 2024b; Chowdhery et al., 2023; dbr, 2025; Liu et al., 2024b). It extends the Transformer architecture by integrating multiple expert networks within the feed-forward network (FFN) component. As illustrated in Figure 2, MoE models dynamically route input tokens to the most relevant experts based on their characteristics. This routing is managed by a trainable gating mechanism that selects the best-suited experts for each token. This architectural innovation enables MoE models to scale in capacity without a proportional increase in inference costs, as only a subset of experts is activated for each input.
2.2. Large-scale LLM Training
@@
Training large language models at scale on tens of thousands of GPUs is a complex system engineering challenge that requires multiple systems techniques. To distribute the training workload, a combination of parallelism strategies such as data, tensor, and pipeline parallelism is necessary (Shoeybi et al., 2019; Rasley et al., 2020; Jiang et al., 2024a), as each approach has limitations that prevent relying on a single method for effective scaling.
@@
Data parallelism uniformly distributes the training data across all devices, with each device replicating the model parameters and optimizer states. To synchronize the parameters after each training iteration, data parallelism performs an all-reduce communication operation. Zero Redundancy Optimizer (ZeRO) (Rajbhandari et al., 2020) improves over data parallelism by distributing model states across all participating devices. ZeRO unfolds across three progressive stages, each designed to increasingly conserve memory, though this comes with the trade-off of elevated communication.
@@
Tensor parallelism distributes compute-intensive tensor operations over multiple devices, enabling parallel computation and significantly accelerating the training process. The specific partitioning strategy and the dependencies among operators within the model dictate that tensor parallelism may necessitate gathering split inputs (all-gather) or merging outputs (reduce-scatter). In LLM training, operators like LayerNorm and Dropout, though less compute-intensive, require substantial activation memory. To tackle this problem, a variant of tensor parallelism known as sequence parallelism (Korthikanti et al., 2023) is proposed, which partitions these operators along the dimension of sequence length. For long-context training, several works (Shoeybi et al., 2019; Jacobs et al., 2023; con, 2025) apply sequence parallelism or tensor parallelism to different operators in self-attention. Figure 3 illustrates the mainstream parallelism strategies for attention, namely tensor, sequence, and context parallelism (TP, SP, and CP), which we analyze in §3.1.
@@
Pipeline parallelism enhances efficiency by dividing model layers into stages that are processed on different devices, enabling pipelined execution. Each batch is split into several micro-batches for this purpose. To minimize pipeline bubbles, various scheduling strategies have been developed, e.g., GPipe (Huang et al., 2019), PipeDream 1F1B (Narayanan et al., 2019) and Interleaved 1F1B (Narayanan et al., 2021), etc. Megatron-LM adopts Interleaved 1F1B pipeline scheduling, further dividing each stage on one device into multiple virtual stages to reduce the pipeline bubble rate.
@@
3. Communication-Efficient Parallelism
@@
Figure 4 shows the design space of parallelism strategies for large-scale MoE training, excluding the outermost data parallelism. We start with inter-node parallelism. Expert parallelism alleviates memory pressure from MoE models’ large parameter size by distributing experts across nodes but incurs per-layer cross-node communication, harming training efficiency. Similarly, tensor parallelism’s high communication overhead makes it more efficient to limit TP to a single node. Following prior work (Jiang et al., 2024a), we adopt pipeline parallelism to distribute model parameters, reduce communication, and overlap communication of different micro-batches.
Prior large-scale MoE training systems, such as Megatron-LM (Shoeybi et al., 2019) and DeepSpeed-MoE (Rajbhandari et al., 2022), incorporate tensor parallelism to scale up training by partitioning the model parameters within the node. However, in our practice, we observe two issues with this approach: (1) TP partitions the expert dimension, which negatively impacts GEMM efficiency; and (2) TP introduces significant communication overhead, which remains constant as the parallelism size increases, eventually causing communication to exceed computation on modern hardware.
@@
3.1. Sequence Parallelism for Attention
Due to the inherent parallelizability of the expert components in MoE models, most prior work on MoE training (Rajbhandari et al., 2022; Li et al., 2023) focuses on optimizing expert parallelism, while data parallelism (DP) is typically applied to the non-MoE components such as attention. However, when scaling up MoE training, this approach proves insufficient due to the activation memory consumption. This issue arises because DP splits the batch dimension both across and within nodes. Compared to other intra-node parallelism strategies shown in Figure 4, applying DP to attention forces each GPU within a node to process one micro-batch simultaneously, increasing the activation size by 8, which often results in out-of-memory issues.
@@ 3.1. Sequence Parallelism for Attention
We adopt sequence parallelism (SP), as proposed in DeepSpeed-Ulysses (Jacobs et al., 2023), to scale MoE training and effectively reduce communication along the critical path. SP is commonly used in long-context training to address memory challenges associated with long inputs. We find it also works well in large-scale MoE training. First, it significantly reduces communication overhead compared to TP, especially when using grouped-query attention (Ainslie et al., 2023). Second, while it introduces some parameter redundancy and increased communication overhead during parameter synchronization, the unique characteristics of MoE models make these trade-offs manageable and acceptable.
@@
where represents the ratio between the number of query heads and that of key-value heads. Assuming the model is trained on an NVIDIA Hopper GPU workstation with an NVLink domain of size 8, the communication latency for sequence parallel attention can be reduced to about one-fourth of that required by tensor-parallel attention.
Data communication & memory overhead. A notable difference between SP and TP attention is how parameters are distributed across devices: TP shards the attention weights, while SP replicates them. This raises the concern about the potential increase in communication overhead for synchronizing gradients and parameters. Counterintuitively, given the intra- and inter-node bandwidth asymmetry and the adoption of hierarchical communication operations in modern communication libraries (NCCL, 2025) as shown in Figure 5 and analyzed in Appendix A.1, although SP attention requires synchronization of more parameters compared to TP attention, the difference in communication overhead is minimal in practical scenarios.
@@
Balanced vs. imbalanced. In addition to the Ulysses-style SP attention, we also explored other forms, including context parallelism (CP) (con, 2025), which partitions all activations along the sequence dimension. CP attention, however, faces workload imbalance due to causal masking in attention, as each token only attends to previous tokens. To mitigate this, we attempted the zigzag strategy by grouping the head and tail partitions of the sequence on the same GPU, although achieving perfect balance remains challenging. Consequently, in large-scale training, the entire training process is often constrained by the most imbalanced data batch. Moreover, this imbalance disturbs the training pipeline, thereby reducing overall training efficiency.
3.2. Expert Parallelism for Feed-forward Network
@@
In practical training, all-to-all communication is less efficient than all-gather and reduce-scatter, as it requires each worker to communicate with all others, whereas all-gather and reduce-scatter follow a ring-based communication pattern with only neighboring workers. As shown in Figure 7, the communication time for these three operations in Mixtral-87B reveals that when top- ¿ 6, the all-gather-based EP implementation is more efficient.
@@
Load balance. A well-known challenge in MoE model training is load balancing across experts (Li et al., 2023; Liu et al., 2024a). To address this, we use auxiliary loss and token dropping to balance the workload across GPUs within each node. Similar to DeepSeek-V2 (Liu et al., 2024a), we treat the experts placed on the same GPU as a group and calculate the balance loss and computational capacity for each device rather than for each individual expert.
4. Communication-computation Overlap
@@
4.1. Inter-operator Overlap
@@
Selective activation rematerialization. The holistic scheduling strategy also helps reduce memory usage without compromising training speed. Compared to dense models with equivalent computational requirements, MoE models exert significantly higher memory pressure during training due to their parameter count being several times larger. In addition to employing ZeRO optimizations (Rajbhandari et al., 2020) to eliminate redundant optimizer states across DP groups, we further optimize memory usage through selective activation rematerialization. This approach reduces activation memory requirements by re-performing computation and communication operators that can be overlapped with other necessary operators.
Figure 8a illustrates the forward pass of a Mixtral (Jiang et al., 2024b) MoE layer and highlights key activations produced during this process. MegaScale-MoE strategically retains activations that are computationally expensive to recompute, while recalculating others generated by memory-intensive operations or communication operations. This minimizes dependencies on backward computation, enabling rematerialization operations to overlap with other computations and communications, avoiding delays in the critical path. For example, as shown in Figure 8b, the backward pass of the GroupedGEMM operator for FC2 requires the activation fc2_in and the gradient of fc2_out (denoted as fc2_out) as inputs. MegaScale-MoE recomputes fc2_in and overlaps this operator with gradient communication (i.e., all-gather for ffn_out). Similarly, ffn_in is obtained through re-performing RMSNorm and all-gather, with these operators hidden within the preceding communication and the FC2 GroupedGEMM, respectively. MegaScale-MoE also places the weighted sum of ffn_out immediately after the SwiGLU (Shazeer, 2020) activation function to eliminate the need to store ffn_out. This reordering ensures computational consistency by avoiding operators that cross non-linear boundaries.
which we have reduced to
MegaScale-MoE reduces the activation memory by while maintaining the same training speed.
4.2. Intra-operator Overlap